### v1

@@ Abstract
Large language model (LLM) inference on resource-constrained edge devices presents a critical challenge for low-latency intelligent systems. State-of-the-art approaches rely on datacenter infrastructure, making local deployment prohibitively expensive. This paper introduces QEIL (Quantifying Edge Intelligence via Inference-time Scaling Laws), a unified framework for efficient LLM inference via principled scaling laws and heterogeneous hardware orchestration across CPU, GPU, and NPU accelerators.
[…]
Comprehensive evaluation across five model families (GPT-2, Granite-350M, Qwen2-0.5B, Llama-3.2-1B, LFM2-2.6B) demonstrates consistent universality: (1) 7–10.5 percentage point improvement in pass@k coverage (66.5%–70.0% vs. 56%–63% baseline), (2) 35.6%–78.2% energy reduction (mean 48.8%), (3) 68% average power reduction to 75–84W (enabling edge thermal budgets), (4) 15.8% average latency improvement, and (5) zero accuracy loss. These improvements hold across the entire model landscape from 125M to 2.6B parameters, establishing that inference-time scaling laws are architecture-agnostic. Heterogeneous orchestration simultaneously achieves superior pass@k coverage, reduced energy consumption, improved latency, and better cost-efficiency (PPP) compared to homogeneous cloud approaches—demonstrating that intelligent local edge execution via hardware-aware task routing is the optimal deployment strategy for energy-constrained AI systems.
@@ 1 Introduction
Efficient inference on heterogeneous edge devices presents one of the most critical challenges in democratizing AI deployment. While training-time scaling laws have been extensively characterized Brown et al. (2024); Hoffmann et al. (2022), the characterization of inference-time scaling laws—particularly across diverse hardware configurations spanning CPUs, GPUs, and NPUs with varying computational capacities—remains fundamentally underexplored. The centralized cloud paradigm, which has dominated LLM deployment, faces mounting infrastructure strain as inference workloads grow exponentially. Recent industry analyses project that global data center capacity must triple by 2030 to meet AI demands Saad-Falcon et al. (2025), yet this centralized approach creates bottlenecks in latency, cost, and environmental sustainability.
Three recent works have made significant strides addressing components of this challenge. First, Brown et al. Brown et al. (2024) demonstrated that inference-time compute scaling through repeated sampling yields log-linear coverage improvements across multiple tasks, suggesting the existence of inference scaling laws. Their work established that coverage—the fraction of problems solved by any generated sample—scales according to an exponentiated power law with sample budget, achieving 4.8× performance gains on SWE-bench by amplifying weaker models through increased sampling. However, their framework focused on single-device, homogeneous execution and did not address the critical challenge of orchestrating inference across heterogeneous hardware with distinct resource constraints (CPU, GPU, NPU) where device-specific cost-efficiency factors differ dramatically.
Second, Saad-Falcon et al. Saad-Falcon et al. (2025) introduced Intelligence Per Watt (IPW) as a unified metric for measuring local inference viability, demonstrating that small local language models can accurately answer 88.7% of single-turn queries with IPW improving 5.3× from 2023-2025. Their work revealed that local-cloud hybrid systems achieve 60-80% energy and cost reductions through intelligent query routing. However, their analysis remained at the query-level routing granularity—treating entire inference requests as atomic units—and did not characterize sub-query-level optimization where different stages of a single inference (e.g., prefill vs. decode) can be routed to different hardware for maximum efficiency.
Third, Asgar et al. Asgar et al. (2025) proposed a comprehensive systems-level framework for agentic AI across heterogeneous infrastructure, introducing MLIR-based representations and cost-aware optimization for decomposed workloads. Their work demonstrated that heterogeneous configurations (e.g., older-generation GPUs paired with newer accelerators) can deliver comparable TCO to homogeneous frontier infrastructure. Critically, they formulated inference execution as a constrained optimization problem over task graphs—a foundational insight for our framework. However, their focus was on agentic workflows with tool calls, memory access, and multiple model invocations, not on characterizing the fundamental scaling law properties of inference on heterogeneous edge devices.
The Gap We Address: Despite these advances, no work has synthesized inference scaling laws with heterogeneous hardware orchestration and principled cost optimization at the inference-time granularity. Specifically, existing work has not established architecture-agnostic theorems quantifying how coverage, energy, cost, and latency scale as functions of model parameters, sample budget and tokens per sample - the core insight required to systematically optimize edge inference systems across diverse hardware. Furthermore, prior work has not provided unified efficiency metrics—beyond IPW—that capture end-to-end energy-coverage trade-offs, nor has it demonstrated empirically that heterogeneous orchestration can simultaneously improve accuracy, reduce latency, and decrease energy consumption compared to homogeneous baselines across diverse model families.
@@
We address these gaps by introducing QEIL (Quantifying Edge Intelligence via Inference-time Scaling Laws), a unified mathematical and systems framework for efficient LLM inference on heterogeneous edge infrastructure spanning CPU, GPU, and NPU devices. Our framework makes three core contributions that build directly on the insights from prior work:
@@
First, we establish five architecture-agnostic theorems quantifying inference-time scaling laws. We prove that coverage scales according to $C(S)=1-\exp(-\alpha N^{\beta}S^{\beta})$ with architecture-agnostic exponent $\beta\approx 0.7$; energy scales sub-linearly with the number of samples as $E(S)=E_{0}\cdot f(Q)\cdot N^{\gamma_{E}}\cdot T\cdot S$ with $\gamma_{E}\approx 0.9$; cost follows similar scaling with device-specific multipliers; and latency scales with parallelism factors. These theorems extend Brown et al.’s empirical observation of log-linear scaling to a complete framework spanning energy, cost, and latency dimensions across heterogeneous devices.
@@
Second, we introduce novel composite efficiency metrics—Energy-Coverage Efficiency (ECE), Intelligence Per Watt (IPW), and Price-Power-Performance (PPP) score—that unify the multi-objective optimization problem. Unlike IPW alone (which captures instantaneous power efficiency), these metrics enable principled comparison across heterogeneous configurations by explicitly modeling the energy-coverage and cost-coverage trade-offs inherent to edge inference.
@@
Third, we propose dynamic heterogeneous orchestration leveraging MLIR-based compilation and cost-aware task placement (inspired by Asgar et al.’s framework) specialized for inference workloads. Our system decomposes inference into granular operations and routes each to the most cost-efficient device—achieving simultaneous improvements in coverage, latency, and energy compared to homogeneous baselines.
@@
Evaluating QEIL across five diverse model families on WikiText—GPT-2 (125M), Granite-350M, Qwen2-0.5B, Llama-3.2-1B, and LFM2-2.6B—across standard (throughput-optimized) and energy-aware (efficiency-optimized) execution paradigms on CPU, GPU, and NPU devices, we achieve: (1) 4.82–5.6× improvement in Intelligence Per Watt (0.718–0.807 vs. 0.130–0.245 baseline), (2) 47.7–78% total energy reduction across models (22,487.8J–212,953.7J vs. baseline), (3) 66.5–70% pass@k coverage versus 56–63% baseline, (4) average 22.5% latency improvement (1.34–1.66ms vs. 1.73–1.91ms), and (5) average 22.9% PPP score improvement (15.49–25.91 vs. 10.44–19.51), all without sacrificing model accuracy. These results empirically validate that heterogeneous edge inference—combined with principled scaling law characterization—fundamentally outperforms homogeneous cloud deployment across diverse model architectures and parameter ranges, demonstrating universality beyond single-model evaluation.
@@
This work makes five primary contributions to edge AI and inference time scaling:
  • We present QEIL, the first framework combining inference-time scaling laws with heterogeneous hardware orchestration across CPU, GPU, and NPU devices. We establish five model-agnostic theorems proving that coverage, energy, cost, and latency follow predictable power-law relationships dependent only on model parameters $N$, sample budget $S$ and tokens per sample TT—not on specific transformer architecture. We validate this universality empirically across five diverse model families (125M–2.6B parameters: GPT-2, Granite-350M, Qwen2-0.5B, Llama-3.2-1B, LFM2-2.6B), proving universal applicability across all model architectures and scales.
@@
  • We demonstrate 4.82–5.6× improvement in Intelligence Per Watt through heterogeneous orchestration across CPU, GPU, and NPU, achieving 66.5–70% pass@k coverage with 47.7–78% energy reduction and 22.5% average latency improvement simultaneously—proving that heterogeneous edge inference surpasses homogeneous cloud infrastructure for realistic workloads across diverse model families. This combination of coverage, energy, and latency improvements is unprecedented in the literature and directly validates our scaling law theorems in practice.
  • We validate empirically that inference-time scaling laws are architecture-agnostic, holding universally across model families (transformers from 125M to 2.6B parameters: GPT-2, Granite, Qwen, Llama, LFM) and parameter counts through mathematical proofs and comprehensive experimental evidence. Our framework scales across diverse model sizes and architectures, demonstrating that our theorems generalize universally across the model landscape and enable practitioners to apply insights to new architectures as they emerge.
  • We introduce a unified heterogeneous computing framework with MLIR-based compilation, cost-aware orchestration, and scaling law validator—enabling reproducible edge intelligence benchmarking validated across five model families and empowering future research to build on these foundations. The framework supports arbitrary transformer architectures, diverse hardware (Intel NPU, Qualcomm NPU, NVIDIA GPU, AMD accelerators), and principled optimization under latency/energy/cost SLAs.
@@ 2.1 Inference-Time Scaling and Repeated Sampling
Recent advances in inference-time compute have emerged as a powerful complement to training-time scaling. Brown et al. (2024) demonstrated that coverage—the fraction of problems solved by any generated sample—scales log-linearly with the number of samples, establishing empirical inference scaling laws across coding, mathematical reasoning, and formal proof domains. Their work showed that on SWE-bench Lite, repeated sampling increased issue resolution from 15.9% with a single attempt to 56% with 250 attempts, validating that inference-compute scaling enables weaker models to surpass stronger single-sample baselines when equipped with sufficient sampling budget. However, their analysis was restricted to uniform sampling budgets and single-device execution without heterogeneous hardware considerations, leaving open the question of how to optimally allocate samples across heterogeneous edge infrastructure with varying energy and computational constraints. Hassid et al. (2024) complemented this by exploring budget reallocation strategies, showing that when constrained by fixed compute budgets (measured in FLOPs), smaller models with more samples can outperform larger models with fewer attempts—a principle foundational to our heterogeneous orchestration strategy.

### v2

@@ Abstract
Large language model (LLM) inference on resource-constrained edge devices presents a fundamental challenge at the intersection of machine learning and systems design. While training-time scaling behavior has been extensively characterized, the scaling behavior of inference under heterogeneous hardware constraints remains poorly understood. In this work, we introduce QEIL (Quantifying Edge Intelligence via Inference-time Scaling Formalisms), a unified framework for characterizing and optimizing inference-time performance across heterogeneous edge devices spanning CPUs, GPUs, and NPUs.
We empirically identify stable scaling relationships governing inference-time coverage, energy, and latency across transformer-based models and heterogeneous hardware configurations. We observe that these relationships generalize across multiple transformer model families spanning 125M–2.6B parameters. Our findings suggest that inference efficiency follows predictable power-law patterns within transformer-based architectures, and that heterogeneous orchestration i.e the intelligent coordination and scheduling of computational workloads across diverse processing units—such as CPUs, GPUs, and NPUs to enable systematic improvements in energy efficiency and task coverage beyond homogeneous execution. Building on these observations, we propose three composite efficiency metrics—Intelligence Per Watt (IPW), Energy-Coverage Efficiency (ECE), and Price-Power-Performance (PPP)—to unify multi-objective optimization across heterogeneous inference configurations.
Critically, QEIL introduces a safety-first agentic orchestration framework for heterogeneous edge environments. Our intelligent agentic orchestrator not only distributes workloads across accelerators from both the same vendor (e.g., Intel CPU paired with Intel NPU and Intel Graphics) and different vendors (e.g., Intel processors combined with NVIDIA GPUs), but also implements comprehensive reliability and AI safety mechanisms—including thermal throttling protection, fault-tolerant execution with graceful degradation, adversarial robustness through input validation, and hardware health monitoring. We adopt a safety-first, capability-second design philosophy, ensuring that model inferencing operates within safe thermal and power envelopes to prevent device damage, even at the cost of reduced peak efficiency. This approach enables building fault-tolerant edge AI systems that prioritize end-user experience and device longevity over raw performance.
Across five model families (GPT-2, Granite-350M, Qwen2-0.5B, Llama-3.2-1B, LFM2-2.6B), QEIL achieves consistent gains in energy efficiency, latency, and coverage without compromising accuracy or system safety. Our results indicate that inference-time scaling and heterogeneous hardware orchestration jointly define a previously underexplored optimization regime for edge AI deployment. These findings suggest that principled inference-time scaling formalisms can complement training-time scaling, offering a systematic framework for designing energy-efficient, reliable, and safe edge intelligence systems.
@@ 1 Introduction
The rapid proliferation of AI agents and edge computing applications has fundamentally transformed the landscape of AI deployment. As AI workloads transition from centralized cloud infrastructure to distributed edge environments, the challenge of efficient inference on resource-constrained heterogeneous devices has emerged as a critical bottleneck for democratizing AI access. Equally important, edge deployment introduces critical safety and reliability concerns: devices must operate within thermal limits to prevent hardware damage, systems must gracefully handle hardware failures, and inference must be robust to adversarial inputs—concerns that are secondary in datacenter environments but paramount for consumer-facing edge applications.
Asgar et al. Asgar et al. (2025) recently presented a seminal systems-level framework for agentic AI workloads across heterogeneous infrastructure, introducing MLIR-based representations and dynamic cost-aware orchestration for datacenter-scale deployments. Their work made several foundational contributions: (1) demonstrating that AI agent workloads can be decomposed into granular components exhibiting distinct sensitivity to hardware resources (TFLOPS, memory bandwidth, network bandwidth), (2) formulating inference execution as a constrained optimization problem over task graphs, and (3) showing that heterogeneous configurations—combining accelerators from different vendors and performance tiers—can deliver comparable or superior TCO to homogeneous frontier infrastructure. Their preliminary finding that H100::Gaudi3 configurations can match B200::B200 performance represents a paradigm shift in infrastructure economics.
However, Asgar et al.’s framework, while transformative for datacenter-scale agentic AI, leaves several critical gaps when applied to resource-constrained edge computing environments:
(1) Edge-Specific Resource Constraints: Datacenter orchestration assumes abundant computational resources, high-bandwidth interconnects (RoCE, InfiniBand), and effectively unlimited power budgets. Edge devices operate under fundamentally different constraints: strict power envelopes (5-85W vs. 300W+ datacenter GPUs), limited memory capacity (8-128GB vs. terabyte-scale), thermal throttling in fanless enclosures, and intermittent network connectivity. Asgar et al.’s optimization framework does not model these edge-specific constraints.
(2) Scaling Relationship Characterization: While Asgar et al. demonstrated TCO benefits through heterogeneous orchestration, they did not establish empirically grounded scaling formalisms quantifying how coverage, energy, and latency scale as functions of model parameters, sample budget, and device characteristics—the foundational insight required to systematically optimize edge inference systems where every joule matters.
(3) Multi-Vendor Edge Heterogeneity: Edge deployments increasingly feature heterogeneous hardware from multiple vendors within a single device (e.g., Intel CPU + Intel NPU + NVIDIA GPU, or Qualcomm NPU + ARM CPU). Asgar et al.’s framework focused on datacenter accelerators (H100, Gaudi3, B200) and did not address the unique challenges of orchestrating across consumer-grade, mixed-vendor edge hardware with vastly different driver stacks, memory hierarchies, and power characteristics.
(4) Distributed Edge Orchestration: Beyond single-device heterogeneity, edge computing increasingly requires distributed orchestration across multiple edge nodes—IoT gateways, mobile edge servers, and embedded devices—each with distinct capabilities. This distributed, resource-constrained paradigm requires fundamentally different optimization approaches than the rack-scale homogeneity assumed in datacenter designs.
(5) Safety, Reliability, and Fault Tolerance: Perhaps most critically, datacenter frameworks assume professionally managed infrastructure with redundant cooling, power backup, and 24/7 monitoring. Edge devices operate in uncontrolled environments—laptops on laps generating heat, mobile devices in direct sunlight, IoT sensors in temperature extremes. A safety-first design philosophy that prioritizes device health over peak performance is essential for consumer-grade edge AI, yet existing frameworks optimize purely for efficiency without considering thermal safety margins, graceful degradation under hardware stress, or adversarial robustness.
Two additional recent works have addressed complementary aspects of this challenge. Brown et al. Brown et al. (2024) observed that inference-time compute scaling through repeated sampling yields log-linear coverage improvements, achieving 4.8× performance gains on SWE-bench by amplifying weaker models through increased sampling. However, their framework focused on single-device, homogeneous execution without addressing heterogeneous hardware orchestration. Saad-Falcon et al. Saad-Falcon et al. (2025) introduced Intelligence Per Watt (IPW) as a unified metric for local inference viability, demonstrating 60-80% energy reductions through intelligent query routing. However, their analysis remained at query-level routing granularity and did not characterize sub-query-level optimization where different inference stages can be distributed across heterogeneous devices.
The Gap We Address: Building directly on Asgar et al.’s foundational insight that heterogeneous orchestration enables cost-efficient AI deployment, we extend their framework to resource-constrained edge computing environments by: (1) establishing empirically grounded scaling formalisms quantifying inference scaling behavior on edge hardware, (2) introducing agentic orchestration capabilities for multi-vendor heterogeneous edge configurations, (3) demonstrating that principled scaling characterization enables systematic optimization under strict power, memory, and latency constraints that define edge deployment scenarios, and (4) implementing a safety-first design philosophy with fault-tolerant execution, thermal protection, and adversarial robustness that ensures reliable end-user experiences on consumer-grade hardware. Our work thus bridges the gap between datacenter-scale heterogeneous AI infrastructure (Asgar et al.) and the resource-constrained, safety-critical reality of edge AI deployment.
@@
We address these gaps by introducing QEIL (Quantifying Edge Intelligence via Inference-time Scaling Formalisms), a unified mathematical and systems framework for efficient LLM inference on heterogeneous edge infrastructure spanning CPU, GPU, and NPU devices from both single and multiple vendors. Our framework makes four core contributions that extend Asgar et al.’s datacenter-focused approach to the edge computing paradigm:
@@
First, we empirically identify five stable scaling formalisms quantifying how inference efficiency scales with fundamental parameters under edge constraints. We observe that coverage scales according to $C(S)=1-\exp(-\alpha N^{\beta_{N}}S^{\beta_{S}})$ with scaling exponents $\beta_{N}\approx 0.7$ and $\beta_{S}\approx 0.7$ that remain consistent across transformer-based architectures; energy scales sub-linearly with model size as $E(S)=E_{0}\cdot f(Q)\cdot N^{\gamma_{E}}\cdot T\cdot S$ with $\gamma_{E}\approx 0.9$; cost follows similar scaling with device-specific multipliers; and latency scales with parallelism factors. These formalisms extend Brown et al.’s empirical observation of log-linear scaling and Asgar et al.’s cost-aware optimization to a complete framework spanning energy, cost, and latency dimensions across heterogeneous edge devices with strict resource constraints.
@@
Second, we introduce novel composite efficiency metrics—Energy-Coverage Efficiency (ECE), Intelligence Per Watt (IPW), and Price-Power-Performance (PPP) score—that unify the multi-objective optimization problem for edge deployment. Unlike IPW alone (which captures instantaneous power efficiency), these metrics enable principled comparison across heterogeneous configurations by explicitly modeling the energy-coverage and cost-coverage trade-offs inherent to battery-powered and thermally-constrained edge inference.
@@
Third, we propose agentic heterogeneous orchestration leveraging MLIR-based compilation and cost-aware task placement (extending Asgar et al.’s datacenter framework) specialized for resource-constrained edge workloads. Our system decomposes inference into granular operations and routes each to the most cost-efficient device—whether from the same vendor (Intel CPU + Intel NPU + Intel Graphics) or different vendors (Intel processors + NVIDIA GPU)—achieving simultaneous improvements in coverage, latency, and energy compared to homogeneous baselines while respecting edge-specific power and memory constraints.
Fourth, we implement a comprehensive safety and reliability framework that treats the heterogeneous orchestrator as an intelligent agentic optimizer responsible not only for performance optimization but also for system health and user safety. This includes: (a) thermal throttling protection that proactively reduces workload intensity when device temperatures approach safe limits, preventing overheating-induced hardware damage; (b) fault-tolerant execution with graceful degradation that automatically redistributes workloads when individual devices fail or become unavailable; (c) adversarial robustness through input validation and anomaly detection that prevents malicious inputs from causing system instability; and (d) hardware health monitoring that tracks device utilization, temperature, and error rates to predict and prevent failures. Our “safety-first, capability-second” philosophy accepts modest efficiency trade-offs (typically 3-5%) to guarantee reliable operation—a critical requirement for consumer-facing edge applications where device damage or system crashes destroy user trust.
@@
Evaluating QEIL across five diverse model families on WikiText—GPT-2 (125M), Granite-350M, Qwen2-0.5B, Llama-3.2-1B, and LFM2-2.6B—across standard (throughput-optimized) and energy-aware (efficiency-optimized) execution paradigms on CPU, GPU, and NPU devices, we observe: (1) 4.82–5.6× improvement in Intelligence Per Watt (0.718–0.807 vs. 0.130–0.245 baseline), (2) 47.7–78% total energy reduction across models (22,487.8J–212,953.7J vs. baseline), (3) 66.5–70% pass@k coverage versus 56–63% baseline, (4) average 22.5% latency improvement (1.34–1.66ms vs. 1.73–1.91ms), (5) average 22.9% PPP score improvement (15.49–25.91 vs. 10.44–19.51), and (6) zero thermal throttling events and 100% fault recovery rate under simulated hardware stress, all without sacrificing model accuracy. These results suggest that heterogeneous edge inference—combined with principled scaling characterization and safety-first design—consistently outperforms homogeneous cloud deployment across diverse model architectures and parameter ranges, indicating generalizability across the transformer model landscape and validating the extension of Asgar et al.’s heterogeneous orchestration paradigm to resource-constrained, safety-critical edge environments.
@@
This work makes six primary contributions to edge AI and inference time scaling:
  • We present QEIL, the first framework combining inference-time scaling formalisms with heterogeneous hardware orchestration across CPU, GPU, and NPU devices, extending Asgar et al.’s datacenter-focused heterogeneous AI framework to resource-constrained edge computing environments. We empirically identify five stable scaling relationships showing that coverage, energy, cost, and latency follow predictable power-law relationships dependent on model parameters $N$, sample budget $S$ and tokens per sample $T$. We validate these formalisms empirically across five diverse transformer model families (125M–2.6B parameters: GPT-2, Granite-350M, Qwen2-0.5B, Llama-3.2-1B, LFM2-2.6B), demonstrating generalizability within the transformer architecture family on our tested hardware platform.
@@
  • We observe 4.82–5.6× improvement in Intelligence Per Watt through agentic heterogeneous orchestration across CPU, GPU, and NPU from both same-vendor and multi-vendor configurations, achieving 66.5–70% pass@k coverage with 47.7–78% energy reduction and 22.5% average latency improvement simultaneously—indicating that heterogeneous edge inference can surpass homogeneous cloud infrastructure for realistic workloads across diverse model families. This combination of coverage, energy, and latency improvements directly validates our scaling formalisms in practice.
  • We find empirically that inference-time scaling relationships exhibit consistent patterns across transformer model families (125M to 2.6B parameters: GPT-2, Granite, Qwen, Llama, LFM), parameter counts, and diverse reasoning benchmarks (WikiText-103, GSM8K, ARC-Challenge) through comprehensive experimental evidence and ablation studies. Our framework demonstrates stable scaling exponents across diverse model sizes and architectures on our tested platform, enabling practitioners to apply these insights within the transformer family.
  • We introduce a unified heterogeneous computing framework with MLIR-based compilation, cost-aware orchestration, and scaling relationship validator—enabling reproducible edge intelligence benchmarking validated across five model families and empowering future research to build on these foundations. The framework supports diverse transformer architectures, diverse hardware from multiple vendors (Intel NPU, Qualcomm NPU, NVIDIA GPU, AMD accelerators), and principled optimization under latency/energy/cost SLAs, demonstrating the practical viability of agentic orchestration in resource-constrained edge environments.
  • We introduce a safety-first reliability framework that implements thermal protection, fault-tolerant execution, adversarial robustness, and hardware health monitoring as first-class concerns in heterogeneous edge orchestration. By treating the orchestrator as an intelligent agentic optimizer responsible for both performance and system health, we demonstrate that “safety-first, capability-second” design achieves robust, reliable operation with minimal efficiency overhead (3-5%), enabling practical deployment on consumer-grade hardware where device longevity and user experience are paramount. This contribution addresses the critical gap between research prototypes and production-ready edge AI systems.
@@ 2.1 Inference-Time Scaling and Repeated Sampling
Recent advances in inference-time compute have emerged as a powerful complement to training-time scaling. Brown et al. (2024) observed that coverage—the fraction of problems solved by any generated sample—scales log-linearly with the number of samples, establishing empirical inference scaling relationships across coding, mathematical reasoning, and formal proof domains. Their work showed that on SWE-bench Lite, repeated sampling increased issue resolution from 15.9% with a single attempt to 56% with 250 attempts, suggesting that inference-compute scaling enables weaker models to surpass stronger single-sample baselines when equipped with sufficient sampling budget. However, their analysis was restricted to uniform sampling budgets and single-device execution without heterogeneous hardware considerations, leaving open the question of how to optimally allocate samples across heterogeneous edge infrastructure with varying energy and computational constraints. Hassid et al. (2024) complemented this by exploring budget reallocation strategies, showing that when constrained by fixed compute budgets (measured in FLOPs), smaller models with more samples can outperform larger models with fewer attempts—a principle foundational to our heterogeneous orchestration strategy.